# TorchGDM: A GPU-Accelerated Python Toolkit for Multi-Scale Electromagnetic Scattering with Automatic Differentiation

### v2

@@ III.1 Package structure
III.2 Main classes and their methods
@@
struct_shifted = struct + [$\Delta x$, $\Delta y$, $\Delta z$].
@@
where r_gpm can be either a number (number of dipole pairs automatically distributed inside the structure), or a list of manually defined positions for the effective dipole pairs. Conversion to a single effective electric and magnetic dipole pair is possible via:
@@ III.4 Postprocessing
1# ------ minimum full example
2import torch
3import matplotlib.pyplot as plt
4import torchgdm as tg
5
6# - vacuum environment
7env = tg.env.EnvHomogeneous3D(env_material=1.0)
8
9# - illumination field(s)
10plane_wave = tg.env.freespace_3d.PlaneWave(
11 e0p=1.0, e0s=0.0
12)
13
14# - discretized structure
15structure = tg.struct3d.StructDiscretizedCubic3D(
16 discretization_config=tg.struct3d.volume.cube(l=8),
17 step=30,
18 materials=tg.materials.MatDatabase("GaN"),
19)
20
21# - define and run simulation
22sim = tg.Simulation(
23 structures=[structure],
24 illumination_fields=[plane_wave],
25 environment=env,
26 wavelengths=torch.linspace(550.0, 950.0, 100),
27)
28sim.plot_structure()
29
30# run the simulation
31sim.run()
32
33# calculate cross sections
34cs_results = sim.get_spectra_crosssections()
35
36# plot
37plt.plot(
38 tg.to_np(cs_results["wavelengths"]),
39 tg.to_np(cs_results["ecs"]),
40)
41plt.show()
III.5 Further tools
@@
IV Examples
IV.1 Minimum Example
IV.2 Benchmarks
Runtime and memory consumption
Figure 6b shows the memory requirement, which is identical for whichever hardware is used. Because so far torchGDM only implements a full solver, the memory requirement increases with $N^{2}$ ($N$ being the number of dipoles). This limits the possible size to around 10,000-15,000 dipoles. In the future we might implement an iterative solver to enable larger simulations.
Comparison with Mie theory
Figure 8: Polarizability extraction accuracy: Mie theory (solid blue lines) is compared to discretization (green, dashed), global polarizability matrix (GPM, orange dashed) and a dipole-pair model (red, dotted). (a) 3D case of a sphere and (b) 2D case of an infinite cylinder. The diameters of the sphere and of the cylinder are both $D=450$ nm. The embedding environment is vacuum, and the material is Gallium Nitride (GaN) in both cases. (a) Cubic discretization with step of 25 nm. (b) Square discretization with step 15 nm. The illumination is a linearly polarized plane wave. In (b) polarization is along the cylinder axis (transverse magnetic, TM). The effective polarizability models are obtained from Mie theory. The 3D GPM model uses 15 dipole pairs, the 2D GPM uses 5 dipole pairs. The single effective dipole approximation (“$\alpha_{\text{eff}}$”) is valid only in the long wavelength region.
Comparison with the T-matrix method
Figure 9: Multi-scattering accuracy: Comparison with T-matrix method. A random arrangement of dielectric spheres ($\epsilon_{\text{sph}}=9$) in vacuum ($\epsilon_{\text{env}}=1$) with radii of $80$ nm, $100$ nm and $120$ nm is laying in the $XY$ plane. The ensemble is illuminated from the left by a plane wave with $\lambda_{0}=800$ nm wavelength. The electric field intensity is calculated at $z_{0}=-250\,$nm below the plane of the sphere centers, using different methods. (a) T-matrix method using “treams”, (b) fully discretized (cubic, step of $20$ nm) and (c) using effective ED, MD dipole pairs. Shown areas are $5\times 5$ µm2. The colormaps are normalized to the same scale in all panels.
The coupled effective dipole description is formally identical to the T-matrix method at dipolar order. Therefore, using small enough spheres described by polarizabilities derived from analytical Mie theory, torchGDM results exactly match Mie theory. On the other hand, GPMs with multiple dipoles per structure are empirical models, obtained by matching the scattered fields outside of the particle and therefore will not exactly match Mie theory or T-Matrix calculations. However, in practice the errors are generally very low in the order of $310^{-}$ 3 or less. Keep in mind that when using volume discretization for polarizability model extraction, the accuracy cannot exceed the fidelity of the discretized simulation. The additional error depends primarily on the number of dipoles in an effective model. Using 10 to 15 dipole pairs is a good rule of thumb to get a very good effective model even for larger structures. A comparison to the T-matrix method is shown in figure 9, further examples comparing 2D and 3D geometries with T-matrix calculations as well as a demonstration how effective models can be obtained from T-matrices are available in the online documentation.
IV.3 Example: Bound States in the Continuum
IV.4 Example: Autograd for resonance search
1# ------ automatic differentiation usage
2# - set require gradients
3wavelengths = torch.as_tensor([700])
4wavelengths.requires_grad = True
5
6# - define and run simulation
7sim = tg.Simulation(
8 ...,
9 wavelengths=wavelengths
10)
11sim.run()
12
13# - calculate *anything*, e.g. the ecs:
14res_ext = tg.postproc.crosssect.ecs(
15 sim, wavelength=wavelengths[0]
16)
17ecs = res_ext["ecs"]
18
19# - gradient: d ecs / d wavelength:
20ecs.backward()
21decs_dwl = wavelengths.grad[0]
IV.5 Example: Huygens metalens optimization
Figure 12: Example of a diffractive lens design by gradient optimization of the positions of silicon cuboids of size $250\times 250\times 125\,$nm3 in the $xy$ plane. (a) Improvement of the field intensity enhancement at the target focus position (0,0, -4000 nm) over 150 iterations of gradient optimization. The insets show the geometries at selected moments during the optimization, indicated by cross markers with the same labels. Inset areas are $6.4\times 6.4\,$µm2. (b) field intensity map of the final design in the focal plane (parallel to $XY$ at $z=-4000$ nm). (c) field intensity map in the $XZ$ plane (parallel to the illumination propagation direction). Both maps show areas of $9\times 9\,$µm2. Illumination is an $x$-polarized normal incident plane wave from the positive towards the negative z-axis, in vacuum, wavelength $\lambda_{0}=775$ nm.
IV.6 Example: Mixed discretization. Quantum emitter within a hollow nanostructure close to many scatterers
1# ------ hybrid mesh / eff. dipole simulation
2# - setup
3env = tg.env.freespace_3d.EnvHomogeneous3D(env_material=1.0)
4wl = 600.0
5e_inc = tg.env.freespace_3d.PlaneWave()
6
7# - discretized structure
8mat = tg.materials.MatConstant(eps=6.0)
9struct_mesh = tg.struct3d.StructDiscretizedCubic3D(
10 tg.struct3d.volume.cube(8), step=15, materials=mat
11)
12
13# - eff. pola. structure
14struct_eff = struct_mesh.convert_to_gpm(
15 r_gpm=5, wavelengths=wl, environment=env
16)
17struct_eff += [150, 450, 0] # move structure
18
19# - create a mixed simulation
20sim_mixed = tg.simulation.Simulation(
21 structures=[struct_mesh, struct_eff],
22 environment=env,
23 illumination_fields=e_inc,
24 wavelengths=wl,
25)
26
27sim_mixed.run()
Listing 3: Simple mixed discretization simulation setup example.
Figure 13: Example of a “mixed discretization” simulation. An optical corral composed of resonant dielectric nano-rods ($\epsilon=10$) contains a plasmonic split-ring resonator (made from silver). The system is in vacuum, illuminated by an electric dipole emitter (red cross) in the gap of the split ring with $\lambda_{0}=500$ nm. (a-c) Total electric field intensity maps $100$ nm below the structure surface. Panel (a) shows the fully discretized simulation. Replacing the dielectric rods by global polarizability matrices and keeping the split-ring in full discretization (b) gives an excellent approximation to the full simulation, with errors in the order of well below ±$\pm 1$% (d). On the other hand, replacing also the split-ring by an effective GPM model (here: 20 dipole pairs), merely gives a qualitative approximation (c) with errors in the order of $40$% (e).
Figure 13 demonstrates this scenario by the example of a local emitter coupled to a plasmonic split-ring resonator, which is in the center of an optical corral composed of resonant dielectric rods.colasdesfrancsOpticalAnalogyElectronic2001 ; chicanneImagingLocalDensity2002 Such a system is typically difficult to describe because of the different involved scales: The emitter’s emission and the plasmonic response require fine discretization, but also the contributions of the optical Corral are essential for the global optical response (Fig. 13a). The T-Matrix method cannot be used because the emitter lies within the circumscribing sphere of the split ring structure. A global polarizability matrix can in principle be used to describe the split-ring antenna with the local light source (Fig. 13c,e), however, optimizing the extraction process such that the emitter field within the gap is correctly modeled would require manual optimization of the GPM extraction procedure. TorchGDM’s automatic GPM fit gives a model that offers only a qualitative agreement (Fig. 13c,e). A mixed discretization however, reproduces the full simulation with errors of the order of ±$\pm 1$% (Fig. 13b,d).
IV.7 Example: Full fields of large structure assemblies via iterative hybrid discretization evaluation
Figure 14: Example of full field evaluation in large, complex structure assemblies with iterative, mixed discretization simulations. (a) Sketch of the procedure, “focusing” on each structure separately and performing a mixed discretization simulation. One simulation per structure is performed, the full fields are eventually combined. The geometry in this example consists of 20 L-shaped silicon structures, each with arm lengths of $300$ nm (long arm) and $200$ nm (short arm). The width and height of the structure are $100$ nm and $125$ nm, respectively. The structures are positioned on a randomly perturbed grid, with random rotation angles. The GPM model consists of 6 effective dipole pairs. (b) statistics of the error of the reconstructed full fields, the standard deviation is below $0.5$%. (c-e) Field intensity maps at $75$ nm below the structure bottom surface. Calculated (c) with a single, fully discretized simulation. (d) with the iterative, mixed discretization procedure. (e) with all structures modelled as GPMs. The zoom insets on the right show an $700\times 800$ nm2 large area, cut through the center plane of the nano-structures on a logarithmic colorscale, demonstrating the faithful reconstruction of near- and internal fields. Nearfields within the circumscribing sphere are also well reproduced with only GPMs. Illumination is an $x$-polarized plane wave incident from the positive towards the negative y-axis in vacuum, wavelength $\lambda_{0}=550$ nm. (f-g) show electric and magnetic field intensity profiles very close ($25$ nm) above the structure surface (evaluated along the dotted vertical line indicated in the field maps).
As a final example, we demonstrate in figure 14 how mixed effective polarizability / discretization simulations can be used to recover full fields from very large simulations. We consider a large assembly of $N$ complex shape particles (here L-shaped silicon corners), that have by themselves a complex optical response including higher order contributions like quadrupoles. We extract an effective GPM model for the L-shape, consisting in six dipole pairs, which approximates accurately the optical response. However, the drawback of the GPM model is, that internal fields cannot be accessed. But we can use TorchGDM’s capability to combine different structure types in a single simulation, in order to recover the full fields. To do so, we run a simulation where the closest neighbors around a considered structure are fully discretized, to obtain the best possible accuracy for close neighbor interactions. Structures farther away than a specific distance (here we arbitrarily chose $1.25\times\lambda_{0}$), are described by an effective polarizability model (see Fig. 14a). Running $N$ simulations, one “centered” on each nano-structure, we finally recover a very accurate approximation ($\lesssim 0.5$ relative error, see Fig. 14b) for the full internal fields of all nanotructures. Those full fields can furthermore be used to calculate the scattered fields of the full ensemble at a higher accuracy than with the GPM approximation alone (see figures 14b-g). This technique allows to recover full fields of otherwise prohibitively large geometries.
V Conclusions
Availability and source code
Online documentation
Acknowledgements.

### v3

@@ II.7 Automatic differentiation limitations
Figure 5: Gallery of the main features of torchGDM. a) Structures can be arbitrarily composed of discretized geometries (here a dielectric prism) and effective models (here small gold nano-rods and a dielectric disc). b) Electric and magnetic near fields. c) Field gradients, either calculated via automatic differentiation or by finite differences (faster). d) Partial, electric and magnetic Local density of states as well as Green’s tensors. e) Rasterscan simulations with scanning illuminations. Here, a focused Gaussian (through NA 1.4, linear $X$-polarization) is raster-scanned across the structure and the scattered far-field intensity is plotted for each beam position. f) Internal fields can be obtained for all discretized structures. g) Radiation patterns (here: cut through XY plane). h) Cross section spectra. i) Exact multipole decomposition (discretized structures only). Here, the multipole decomposition (up to quadrupoles) of the contribution of the discretized Si3N4 triangular structure to the scattering cross section is plotted. j) Near-field enhancement spectra (here $100$ nm below the center of the triangle). All 2D plots show an area of $800\times 800$ nm2. Except for the LDOS and rasterscan examples, the illumination is a plane wave in the structure plane, coming from the positive $Y$ axis, with linear polarization along $X$. Single wavelength results are calculated at $\lambda_{0}=550\,$nm. The host medium is vacuum.
II.8 Package structure
We limit the presentation here to the basic package structure and the main classes. The most relevant elements of the interface are listed and discussed in the appendix A. For a full technical documentation we refer the reader to the online documentation at https://gitlab.com/wiechapeter/torchgdm.
Subpackages:
@@ II.8 Package structure
Main classes:
There are two main classes, defined at the top level of torchGDM. The first top-level class “tg.Simulation” provides an API for all simulation related postprocessing and visualization functionality of torchGDM. Complex fields (near- and far-field) are generally returned as instances of the second top-level class “tg.Field”, which provides an API for various post-processing and visualizations related to the electromagnetic fields.
Results:
While electromagnetic fields are returned as instances of the tg.Field class, other results are usually returned in dictionaries, for which torchGDM tries to maintain a consistent nomenclature. Observables related to the scattered, total and incident field are returned in dictionary keys carrying the suffixes “sca”, “tot” and “inc”, respectively. Scattering, extinction and absorption cross widths (2D) and cross sections (3D) are indicated by suffixes “scs”, “ecs”, “acs”.
API:
Most functions are accessible through an object oriented interface from the torchGDM classes, but also through a functional interface in the different subpackages. For a detailed discussion of all submodules, functions and class methods, please see Appendix A or the online API documentation (https://gitlab.com/wiechapeter/torchgdm).
1# ------ minimum full example
2import torch
3import matplotlib.pyplot as plt
4import torchgdm as tg
5
6# - vacuum environment
7env = tg.env.EnvHomogeneous3D(env_material=1.0)
8
9# - illumination field(s)
10plane_wave = tg.env.freespace_3d.PlaneWave(
11 e0p=1.0, e0s=0.0
12)
13
14# - discretized structure
15structure = tg.struct3d.StructDiscretizedCubic3D(
16 discretization_config=tg.struct3d.volume.cube(l=8),
17 step=30,
18 materials=tg.materials.MatDatabase("GaN"),
19)
20
21# - define and run simulation
22sim = tg.Simulation(
23 structures=[structure],
24 illumination_fields=[plane_wave],
25 environment=env,
26 wavelengths=torch.linspace(550.0, 950.0, 100),
27)
28sim.plot_structure()
29
30# run the simulation
31sim.run()
32
33# calculate cross sections
34cs_results = sim.get_spectra_crosssections()
35
36# plot
37plt.plot(
38 tg.to_np(cs_results["wavelengths"]),
39 tg.to_np(cs_results["ecs"]),
40)
41plt.show()
III Examples
III.1 Capabilities
To give an overview of what can be done with torchGDM, some of the core capabilities are summarized in figure 5. An extensive list of all implemented observables is given in the Appendix or in the online documentation. All shown calculations are fully compatible with PyTorch automatic differentiation.
So far, only homogeneous media are supported, but torchGDM can be easily extended without modification of the core code through custom environment classes. An example showing how to implement a host medium with a dielectric interface through a mirror dipole approximation is given in the online documentation.
III.2 Minimum Example
III.3 Benchmarks
Runtime and memory consumption
Figure 7b shows the memory requirement, which is identical for whichever hardware is used. Because so far torchGDM only implements a full solver, the memory requirement increases with $N^{2}$ (NN being the number of dipoles). This limits the possible size to around 10,000-15,000 dipoles. In the future we might implement an iterative solver to enable larger simulations.
Accuracy of effective dipole models
Figure 9: Demonstration of the accuracy of the effective models in comparison with the full discretization of a complex 2D nanostructure. (a) Sketch of the geometry and the effective model (black markers). Red and green markers indicate, respectively, the source and probe locations used for the effective model extraction. The full discretization is composed of roughly 3700 dipoles on a 2D square mesh with a step of $14.2$ nm. The effective GPM structure is a global polarizability matrix (GPM) based on 20 dipole pairs. The structure is placed in vacuum and is made of a lossless dielectric with $\epsilon=9$. It is illuminated by a line dipole oscillating along the $Y$ axis, placed at the location indicated by the blue marker. (b) Radiation pattern of full discretization (solid blue line) vs effective dipole model (dashed orange line). (c) Total electric nearfield intensity calculated from the discretized structure. (d) Same as (c) but calculated using the effective model. White markers indicate the GPM effective dipole locations.
Figure 9 demonstrates the accuracy of the global polarizability matrix effective models for a complicated 2D scatterer, using a difficult illumination. The structure is shown in figure 9a, it is made of a lossless dielectric with $\epsilon=9$, and is placed in vacuum. The discretized model consists of a bit more than 3700 dipoles, the equivalent effective model is made from 20 dipole pairs, corresponding to a reduction of roughly $\times 100$. The illumination is a line dipole along $Y$, placed within a notch of the geometry. The location of the local light source is well within the circumscribing circle of the particle, thus this situation cannot be modelled with conventional T-Matrix in vector spherical harmonics expansion. Figure 9b shows an excellent agreement between the farfield scattering patterns of the discretized model vs the global polarizability matrix. Figure 9c-d compare the electric nearfield intensity maps of the two cases. Outside of the extraction probe and source locations (green and red dots in Fig. 9a), the near-fields are very accurately modelled by the GPM (also within the circumscribing circle).
Comparison with Mie theory
Figure 10: Polarizability extraction accuracy: Mie theory (solid blue lines) is compared to discretization (green, dashed), global polarizability matrix (GPM, orange dashed) and a dipole-pair model (red, dotted). (a) 3D case of a sphere and (b) 2D case of an infinite cylinder. The diameters of the sphere and of the cylinder are both $D=450$ nm. The embedding environment is vacuum, and the material is Gallium Nitride (GaN) in both cases. (a) Cubic discretization with step of 25 nm. (b) Square discretization with step 15 nm. The illumination is a linearly polarized plane wave. In (b) polarization is along the cylinder axis (transverse magnetic, TM). The effective polarizability models are obtained from Mie theory. The 3D GPM model uses 15 dipole pairs, the 2D GPM uses 5 dipole pairs. The single effective dipole approximation ($\alpha_{\text{eff}}$”) is valid only in the long wavelength region.
Comparison with the T-matrix method
Figure 11: Multi-scattering accuracy: Comparison with T-matrix method. A random arrangement of dielectric spheres ($\epsilon_{\text{sph}}=9$) in vacuum ($\epsilon_{\text{env}}=1$) with radii of $80$ nm, $100$ nm and $120$ nm is laying in the XYXY plane. The ensemble is illuminated from the left by a plane wave with $\lambda_{0}=800$ nm wavelength. The electric field intensity is calculated at $z_{0}=-250\,$nm below the plane of the sphere centers, using different methods. (a) T-matrix method using “treams”, (b) fully discretized (cubic, step of 2020 nm) and (c) using effective ED, MD dipole pairs. Shown areas are $5\times 5$ µm2. The colormaps are normalized to the same scale in all panels.
The coupled effective dipole description is formally identical to the T-matrix method at dipolar order. Therefore, using small enough spheres described by polarizabilities derived from analytical Mie theory, torchGDM results exactly match Mie theory. On the other hand, GPMs with multiple dipoles per structure are empirical models, obtained by matching the scattered fields outside of the particle and therefore will not exactly match Mie theory or T-Matrix calculations. However, in practice the errors are generally very low in the order of $10^{-}3$ or less. Keep in mind that when using volume discretization for polarizability model extraction, the accuracy cannot exceed the fidelity of the discretized simulation. The additional error depends primarily on the number of dipoles in an effective model. Using 10 to 15 dipole pairs is a good rule of thumb to get a very good effective model even for larger structures. A comparison to the T-matrix method is shown in figure 11, further examples comparing 2D and 3D geometries with T-matrix calculations as well as a demonstration how effective models can be obtained from T-matrices are available in the online documentation.
III.4 Example: Bound States in the Continuum
III.5 Example: Autograd for resonance search
1# ------ automatic differentiation usage
2# - set require gradients
3wavelengths = torch.as_tensor([700])
4wavelengths.requires_grad = True
5
6# - define and run simulation
7sim = tg.Simulation(
8 ...,
9 wavelengths=wavelengths
10)
11sim.run()
12
13# - calculate *anything*, e.g. the ecs:
14res_ext = tg.postproc.crosssect.ecs(
15 sim, wavelength=wavelengths[0]
16)
17ecs = res_ext["ecs"]
18
19# - gradient: d ecs / d wavelength:
20ecs.backward()
21decs_dwl = wavelengths.grad[0]
III.6 Example: Huygens metalens optimization
Figure 14: Example of a diffractive lens design by gradient optimization of the positions of silicon cuboids of size $250\times 250\times 125\,$nm3 in the xyxy plane. (a) Improvement of the field intensity enhancement at the target focus position ($0,$ -4000 nm) over 150 iterations of gradient optimization. The insets show the geometries at selected moments during the optimization, indicated by cross markers with the same labels. Inset areas are $6.4\times 6.4\,$µm2. (b) field intensity map of the final design in the focal plane (parallel to XYXY at z=4000z=-4000 nm). (c) field intensity map in the XZXZ plane (parallel to the illumination propagation direction). Both maps show areas of $9\times 9\,$µm2. Illumination is an $x$-polarized normal incident plane wave from the positive towards the negative z-axis, in vacuum, wavelength $\lambda_{0}=775$ nm.
III.7 Example: Mixed discretization. Quantum emitter within a hollow nanostructure close to many scatterers
1# ------ hybrid mesh / eff. dipole simulation
2# - setup
3env = tg.env.freespace_3d.EnvHomogeneous3D(env_material=1.0)
4wl = 600.0
5e_inc = tg.env.freespace_3d.PlaneWave()
6
7# - discretized structure
8mat = tg.materials.MatConstant(eps=6.0)
9struct_mesh = tg.struct3d.StructDiscretizedCubic3D(
10 tg.struct3d.volume.cube(8), step=15, materials=mat
11)
12
13# - eff. pola. structure
14struct_eff = struct_mesh.convert_to_gpm(
15 r_gpm=5, wavelengths=wl, environment=env
16)
17struct_eff += [150, 450, 0] # move structure
18
19# - create a mixed simulation
20sim_mixed = tg.simulation.Simulation(
21 structures=[struct_mesh, struct_eff],
22 environment=env,
23 illumination_fields=e_inc,
24 wavelengths=wl,
25)
26
27sim_mixed.run()
Listing 3: Simple mixed discretization simulation setup example.
Figure 15: Example of a “mixed discretization” simulation. An optical corral composed of resonant dielectric nano-rods ($\epsilon=10$) contains a plasmonic split-ring resonator (made from silver). The system is in vacuum, illuminated by an electric dipole emitter (red cross) in the gap of the split ring with $\lambda_{0}=500$ nm. (a-c) Total electric field intensity maps $100$ nm below the structure surface. Panel (a) shows the fully discretized simulation. Replacing the dielectric rods by global polarizability matrices and keeping the split-ring in full discretization (b) gives an excellent approximation to the full simulation, with errors well below $\pm 1$% (d). On the other hand, replacing also the split-ring by an effective GPM model (here: 20 dipole pairs), merely gives a qualitative approximation (c) with errors up to 4040% (e).
Figure 15 demonstrates this scenario by the example of a local emitter coupled to a plasmonic split-ring resonator, which is in the center of an optical corral composed of resonant dielectric rods.colasdesfrancsOpticalAnalogyElectronic2001 ; chicanneImagingLocalDensity2002 Such a system is typically difficult to describe because of the different involved scales: The emitter’s emission and the plasmonic response require fine discretization, but also the contributions of the optical Corral are essential for the global optical response (Fig. 15a). The T-Matrix method cannot be used because the emitter lies within the circumscribing sphere of the split ring structure. A global polarizability matrix can in principle be used to describe the split-ring antenna with the local light source (Fig. 15c,e), however, optimizing the extraction process such that the emitter field within the gap is correctly modeled would require manual optimization of the GPM extraction procedure. TorchGDM’s automatic GPM fit gives a model that offers only a qualitative agreement (Fig. 15c,e). A mixed discretization however, reproduces the full simulation with errors of the order of $\pm 1$% (Fig. 15b,d).
III.8 Example: Full fields of large structure assemblies via iterative hybrid discretization evaluation
Figure 16: Example of full field evaluation in large, complex structure assemblies with iterative, mixed discretization simulations. (a) Sketch of the procedure, “focusing” on each structure separately and performing a mixed discretization simulation. One simulation per structure is performed, the full fields are eventually combined. The geometry in this example consists of 20 L-shaped silicon structures, each with arm lengths of $300$ nm (long arm) and $200$ nm (short arm). The width and height of the structure are $100$ nm and $125$ nm, respectively. The structures are positioned on a randomly perturbed grid, with random rotation angles. The GPM model consists of 6 effective dipole pairs. (b) statistics of the error of the reconstructed full fields, the standard deviation is below 0.50.5%. (c-e) Field intensity maps at $75$ nm below the structure bottom surface. Calculated (c) with a single, fully discretized simulation. (d) with the iterative, mixed discretization procedure. (e) with all structures modelled as GPMs. The zoom insets on the right show an $700\times 800$ nm2 large area, cut through the center plane of the nano-structures on a logarithmic colorscale, demonstrating the faithful reconstruction of near- and internal fields. Nearfields within the circumscribing sphere are also well reproduced with only GPMs. Illumination is an $x$-polarized plane wave incident from the positive towards the negative y-axis in vacuum, wavelength $\lambda_{0}=550$ nm. (f-g) show electric and magnetic field intensity profiles very close ($25$ nm) above the structure surface (evaluated along the dotted vertical line indicated in the field maps).
As a final example, we demonstrate in figure 16 how mixed effective polarizability / discretization simulations can be used to recover full fields from very large simulations. We consider a large assembly of $N$ complex shape particles (here L-shaped silicon corners), that have by themselves a complex optical response including higher order contributions like quadrupoles. We extract an effective GPM model for the L-shape, consisting in six dipole pairs, which approximates accurately the optical response. However, the drawback of the GPM model is, that internal fields cannot be accessed. But we can use TorchGDM’s capability to combine different structure types in a single simulation, in order to recover the full fields. To do so, we run a simulation where the closest neighbors around a considered structure are fully discretized, to obtain the best possible accuracy for close neighbor interactions. Structures farther away than a specific distance (here we arbitrarily chose $1.25\times\lambda_{0}$), are described by an effective polarizability model (see Fig. 16a). Running $N$ simulations, one “centered” on each nano-structure, we finally recover a very accurate approximation ($\lesssim 0.5$ % relative error, see Fig. 16b) for the full internal fields of all nanotructures. Those full fields can furthermore be used to calculate the scattered fields of the full ensemble at a higher accuracy than with the GPM approximation alone (see figures 16b-g). This technique allows to recover full fields of otherwise prohibitively large geometries.
IV Conclusions
Availability and source code
Online documentation
Acknowledgements.
APPENDIX
A API Details
We present in this appendix the most relevant elements of the torchGDM interface. For a full technical documentation we refer the reader to the online documentation.
1 Main classes and their methods
@@
struct_shifted = struct + $\Delta x$, $\Delta y$, $\Delta z$].
@@
where r_gpm can be either a number (number of dipole pairs automatically distributed inside the structure), or a list of manually defined positions for the effective dipole pairs. Note that if a number is given for r_gpm, the extraction is not autodiff compatible because and external clustering algorithm will be used. Conversion to a single effective electric and magnetic dipole pair is possible via:
@@ 3 Postprocessing
4 Further tools